The product of two consecutive numbers is 72. Use a quadratic equation to find the two sets of numbers.
So, (x - 8)(x-9) = 0
x^2 + -17 + 72 = 0
ax^2 + bx + c, where a = 1, b = -17 and c = 72.
"Plug" into x = b+-sqrt4ac/2a
I found online that:
-17 + sqrt(1)
-17 + 1 = -16/2 = -8
-17 - 1 = -18/2 = -9
Aren't you meant to plug in all the numbers in the equation? Why am I getting decimal numbers?

Found 2 solutions by phoihe001, Alan3354:
Answer by phoihe001(34)
Answer by Alan3354(69443)

The product of two consecutive numbers is 72. Use a quadratic equation to find the two sets of numbers.
So, (x - 8)(x-9) = 0 ******* You started by knowing the answer.
x^2 + -17 + 72 = 0
ax^2 + bx + c, where a = 1, b = -17 and c = 72.
"Plug" into x = b+-sqrt4ac/2a
I found online that:
-17 + sqrt(1)
-17 + 1 = -16/2 = -8
-17 - 1 = -18/2 = -9
Aren't you meant to plug in all the numbers in the equation? Why am I getting decimal numbers? What decimal numbers?